OpenStudy (anonymous):

1 2/3 divided by 6, and show work.

OpenStudy (anonymous):

1 & 2/3 = 8/3 8/3 : 6= 8/3 * 1/6=8/(3*6)=8/18=4/9

OpenStudy (gw2011):

1 2/3 = 5/3 (5/3)/6 = (5/3) (1/6) = 5/18